# Addressing the sign-problem in Euclidean path integrals with radial basis function neural networks

### v2

@@ I Introduction
First, the RBF method is briefly described in parallel with the description of the underlying system (the interacting complex scalar field) that it will be applied to. Then, after introducing the model, the method is used to calculate the critical chemical potential that corresponds to Bose condensation at finite densities. Finally, the silver blaze region is addressed through the determination of the number density at small chemical potentials. At the end, the future possibilities and extensions of the RBF model are briefly discussed.
@@ II Radial basis function expansion of the path integral
$S_{1,x}(m,\lambda)=\frac{m^{2}}{2}\sum_{a=1,2}\phi_{a,x}^{2}+\frac{\lambda}{4}\Big(\sum_{a=1,2}\phi_{a,x}\Big)^{2}.$ (4)
@@
Regarding the RBF network, a few important notes have to be made here. First, the width of the kernels ($A$) can be kept constant, as it is always possible to train the network to describe a nonlinear functional form by choosing some corresponding centers, weights, and a possible scaling of the fields. Second, the centers ($c_{k}$) have to be chosen carefully to satisfy a few criteria, which will be defined more precisely later. In general, a good parametrization should include a symmetric center distribution (around 0) that covers the range of the function, with the addition that the RBF approximation should decay very fast outside the range where $e^{-S_{1,x}}$ does not have dominant contributions. The former condition comes from the momentum space behavior of the RBF expansion, while the latter makes sure that one does not have additional contributions when integrating out the path integral.
@@
To examine the differences that the $U$ transformation makes in the results, we set $a_{k}=1$ $(\forall k)$ and compare the general expression in Eq. 11 with the same form, but instead of $(U^{T}\hat{c}_{a,k})$, we use $\hat{c}_{a,k_{p}}$ in the 2nd line. We will call the latter $Z_{U=0}$, and the former $Z_{U}$, and compare $\ln Z_{U=0}$ and $\ln Z_{U}$ through their averaged relative errors for different $c_{k}$ distributions defined as:
$E_{R}(N_{c})=\frac{1}{M}\sum_{i=1}^{M}\frac{|\ln Z_{U,i}(N_{c})-\ln Z_{U=0,i}(N_{c})|}{|\ln Z_{U,i}(N_{c})|},$ (13)
@@
In these simulations we have chosen $K=20$ centers randomly from a uniform distribution between the interval $U[a,b]$, and calculated the relative errors for fixed $N_{c}$ combinations. A more detailed description of the error analysis for the free scalar field case can be found in [21]. The results for the case of the complex scalar fields at finite chemical potentials can be seen in Fig. 1 for different $c_{k}\in U[a,b]$ uniformly distributed centers, $A$ widths, and $\mu$ chemical potentials.
Figure 1: Dependence of the relative difference between $\ln Z_{U}$, and $\ln Z_{U=0}$ on the number of combinations $N_{c}$ included for different RBF parametrizations and chemical potentials.
According to the results, by using symmetric $c_{k}$ distributions around 0 and some corresponding $A$ parameters, it seems possible to estimate $\ln Z$ with very good accuracy, using Eq. 9, with $U$ set to 1. It is also evident that the error converges to a fixed value after a few hundred samples, and while the error depends on the distribution of the centers, their intervals, and the $A$ and $\mu$ parameters, it stays below a few percent if one chooses symmetric centers around 0.
@@ III Results
Using the RBF approximation of the generating functional makes it possible to calculate observables in a very convenient way. In general, due to the Gaussian nature of the expressions, the finite chemical potential should not pose any further problems. Here, we will be interested in two phenomena related to the relativistic Bose gas at finite densities. First, we would like to address the phase transition at some critical chemical potential and determine $\mu_{c}$ by using the RBF approximation. Secondly, we will address the silver-blaze problem, that is, the $\mu$-independent behavior of the observables at zero temperatures (i.e., large $N_{t}$) when $\mu<\mu_{c}$.
@@
$S_{1,x}=\frac{m^{2}}{2}\!\!\!\sum_{a=1,2}\phi_{a,x}^{2}+\frac{\lambda}{4}\Big(\sum_{a=1,2}\phi_{a,x}\Big)^{2}\!\!+\!\!\sum_{a=1,2}J_{a}\phi_{a,x}$ (15)
@@
where $Z^{rbf}$ is the RBF approximated partition function. The $\mu$ independence is the consequence of severe cancellations due to the imaginary parts in the action, thus, in the phase quenched theory (i.e., when $\mathcal{B}_{p}=0$ in Eq. II) we do not expect this behavior. By having the RBF approximated partition function, one could derive the closed-form expression for the expectation value of $\langle n\rangle$, or one could take the numerical derivatives by, e.g., a simple forward difference scheme. In this work we choose the latter with $\Delta\mu=0.001$. The calculations are done up until the critical chemical potential at a lattice size of $N=10$, and the results can be followed in Fig. 4, where the phase quenched values (i.e., when $\mathcal{B}_{p}=0$) are also shown for comparison.
@@
IV Conclusions
@@
The RBF method is able to provide results in a very fast manner, e.g., the calculation regarding the silver blaze problem in Fig. 4 on a $10^{4}$ lattice took only $72$ seconds on a standard notebook without any optimization or using any GPU or special hardware. Due to the generality of the model, there is a possibility to extend it to include fermionic fields and non-abelian gauge theories and address problems in QCD at finite densities as well. On the technical side, due to its versatility, the RBF model could also be extended or formulated in other ways that could be more suitable to specific problems. One alternative would be to also include the hopping terms in the RBF expansion, thus approximating the nearest neighbor interactions by a separable form. This could make the diagonalization step and the error estimation for different center distributions unnecessary, as the complexity will greatly reduce in coordinate space as well. Other alternatives of the RBF model would also be possible, where the effectiveness of each model generally depends on the specific system being analyzed.
V acknowledgments
@@ Acknowledgements.
This work was supported by the Korea National Research Foundation under Grant No. 2023R1A2C300302311 and 2023K2A9A1A0609492411.

### v3

@@ I Introduction
First, the RBF method is briefly described in parallel with the description of the underlying system (the interacting complex scalar field) that it will be applied to. Then, after introducing the model, the method is used to calculate the critical chemical potential that corresponds to Bose condensation at finite densities. Finally, the silver blaze region is addressed through the determination of the number density at small chemical potentials. At the end, numerical considerations, time complexity analysis, and future possibilities and extensions of the RBF model are briefly discussed.
@@ II Radial basis function expansion of the path integral
$S_{1,x}(m,\lambda)=\frac{m^{2}}{2}\sum_{a=1,2}\phi_{a,x}^{2}+\frac{\lambda}{4}\Big(\sum_{a=1,2}\phi_{a,x}^{2}\Big)^{2}.$ (4)
@@
According to Eq. 5, the corresponding RBF network is a multiple-input-single-output (MISO) system that is defined at each lattice with inputs $\phi_{1,x}$ and $\phi_{2,x}$ and an output that aims to approximate $e^{-S_{1,x}(m,\lambda)}$. Its general structure can be followed in Fig. 1, where the first layer consists of the inputs, the middle (hidden) layer consists of $K$ number of nonlinear radial activation functions $F_{1}$, $F_{2}$, …, $F_{K}$, while the final (linear) layer corresponds to a weighted sum of the outputs of the hidden layer with weight factors $a_{1}$, $a_{2}$, …, $a_{K}$.
Figure 1: General structure of the RBF network with two inputs $\phi_{1,x}$, $\phi_{2,x}$, and $K$ activation functions in the hidden layer. The output corresponds to the weighted sum of the $F_{1},F_{2},...,F_{K}$ Gaussian activation functions that have a constant AA width parameter and k-dependent $c_{1,k}$, $c_{2,k}$ centers. The $S_{c}$ weights acts as a scaling parameter of the fields, and is set to $S_{c}=1$.
The $S_{c}$ parameters that are shown between the input and the hidden layers are correspond to a possible scaling of the inputs as $\phi\rightarrow S_{c}\phi$, which could be important if one would want to constraint the inputs to a predefined interval. In this work $S_{c}$ is always set to $1$, thus it will not be shown explicitly in the equations. In the hidden layer each nonlinear activation function is a Gaussian that is parametrized by its centers $(c_{1,k},c_{2,k})$, and its width $(A)$.
Regarding the RBF network, a few important notes have to be made here. First, the width of the kernels ($A$) can be kept constant, as it is always possible to train the network to describe a nonlinear functional form by choosing some corresponding centers, weights, and a possible scaling of the fields. Second, the centers $(c_{1,k},c_{2,k})$ have to be chosen carefully to satisfy a few criteria, which will be defined more precisely later. In general, a good parametrization should include a symmetric center distribution (around zero) that covers the range of the function, with the addition that the RBF approximation should decay very fast outside the range where $e^{-S_{1,x}}$ does not have dominant contributions. The former condition comes from the momentum space behavior of the RBF expansion, while the latter makes sure that one does not have additional contributions when integrating out the path integral.
@@
Next, a simple example will be shown in the case when the mass is set to $m=1$, and the coupling is defined as $\lambda=2$ to show one possible training procedure of the RBF network. Putting these values into $S_{1,x}(m,\lambda)$, the RBF network can be trained to approximate its values on a finite operating range of $\phi_{1,x}$ and $\phi_{2,x}$. Due to the exponential damping in $e^{-S_{1,x}(m,\lambda)}$, it is an easy task to determine the necessary range for the centers and widths of the Gaussian basis functions. As $e^{-S_{1,x}(m,\lambda)}$ with this parametrization only gives non-negligible values in the range between $\phi_{1,x}\in[-2,2]$ and $\phi_{2,x}\in[-2,2]$ the centers are set to cover a grid between $(c_{1,k},c_{2,k})\in[-1,1]\times[-1,1]$ with $K=10^{2}$ centers distributed at equal $\Delta c=(1+1)/(10-1)$ distances in both directions. The width of each Gaussian is set to $A=3$, which is a good choice for these parameterizations. In general, the first step of the training procedure is always the determination of the centers and the widths, and their values depend on the approximable functions. If the function is heavily oscillating, a larger $A$ value could be necessary to be able to capture the high-frequency parts as well. In practice the number of kernels, the distribution of the centers, and the value of the width parameter need to be determined in a way so that the training of the $a_{k}$ weights can converge to a satisfying value, and the approximable function can be described with good accuracy. To determine the $a_{k}$ weight parameters, $N_{T}=1000$ training points are generated randomly in the range of $(\phi_{1,x},\phi_{2,x})\in[-2.5,2.5]\times[-2.5,2.5]$ and the optimization problem is solved in the least squares sense by first constructing the coefficient matrix $H(c_{k},A)\in\mathbb{R}^{N_{T}\times K}$ and then solving the optimization problem described as follows:
$\hat{a}=\arg\min_{a}\;\|H(c_{k},A)a-y\|^{2},$ (6)
where $a=(a_{1},a_{2},...,a_{K})$ is the sought weight parameter vector, and $y=(y_{1},y_{2},...,y_{N_{T}})$ is the generated output vector, while the $H(c_{k},A)$ coefficient matrix depends on the width parameter $A$ and on the centers $c_{k}=(c_{1,k},c_{2,k})$. In general this optimization problem can be solved in many ways, e.g., by gradient descent algorithms, singular value decomposition (SVD), etc. In this case the SVD method is used with a small $r_{c}=0.001$ regularization parameter that helps prevent instabilities in the numerical method by truncating the very small singular values. In general, as the coefficient matrix $H$ depends on the $A$ and $c_{k}$ parameters, the full optimization procedure should also consist of the determination of the number of kernels $K$, the widths of the Gaussians $A$, and their centers $c_{k}$. The easiest way to do this is to generate many different configurations, do the optimization shown in Eq. 6, and then choose the best model that has the lowest mean squared error for a randomized test set. A more sophisticated method to address the complexity of the model is to use, e.g., the Akaike Information Criterion (AIC) [4], which is able to compare the different models in a more standardized way. The comparison between the true values and the RBF approximation for the given problem ($m=1$, $\lambda=2$) with the parameters ($A=3$, $c_{k}\in[-1,1]\times[-1,1]$, $K=10^{2}$) can be seen in Fig. 2 using a test set in the interval of $(\phi_{1},\phi_{2})\in[-2,2]\times[-2,2]$ on a uniform grid with a grid resolution of $\Delta=0.02$.
[…]
@@
To examine the differences that the $U$ transformation makes in the results, we set $a_{k}=1$ $(\forall k)$ and compare the general expression in Eq. 12 with the same form, but instead of $(U^{T}\hat{c}_{a,k})_{p}$, we use $c_{a,k_{p}}$ in the 2nd line. We will call the latter $Z_{0}$, and the former $Z_{U}$, and compare $\ln Z_{0}$ and $\ln Z_{U}$ through their averaged relative errors for different $c_{k}$ distributions defined as:
$E_{R}(N_{c})=\frac{1}{M}\sum_{i=1}^{M}\frac{|\ln Z_{U,i}(N_{c})-\ln Z_{0,i}(N_{c})|}{|\ln Z_{U,i}(N_{c})|},$ (14)
@@
In these simulations we have chosen $K=20$ centers randomly from a uniform distribution between the interval $[a,b]$, and calculated the relative errors for fixed $N_{c}$ combinations. A more detailed description of the error analysis for the free scalar field case can be found in [8]. The results for the case of the complex scalar fields at finite chemical potentials can be seen in Fig. 3 for different $c_{k}\sim Uniform[a,b]$ uniformly distributed centers, $A$ widths, and $\mu$ chemical potentials.
Figure 3: Dependence of the relative difference between $\ln Z_{U}$, and $\ln Z_{0}$ on the number of combinations $N_{c}$ included for different RBF parametrizations and chemical potentials.
According to the results, by using symmetric $c_{k}$ distributions around 0 and some corresponding $A$ parameters, it seems possible to estimate $\ln Z$ with very good accuracy, using a factorized form in momentum space shown in Eq. 11. It is also evident that the error converges to a fixed value after a few hundred samples, and while the error depends on the distribution of the centers, their intervals, and the $A$ and $\mu$ parameters, it stays below a few percent if one chooses symmetric centers around 0.
@@ III Results
Using the RBF approximation of the generating functional makes it possible to calculate observables in a very convenient way. In general, due to the Gaussian nature of the expressions, the finite chemical potential should not pose any further problems. Here, we will be interested in two phenomena related to the relativistic Bose gas at finite densities. First, we would like to address the phase transition at some critical chemical potential and determine $\mu_{c}$ by using the RBF approximation. Secondly, we will address the silver-blaze problem, that is, the $\mu$-independent behavior of the observables at zero temperatures (i.e., large $N_{t}$) when $\mu<\mu_{c}$. In both cases the system parameters are set to $m=1$, and $\lambda=1$.
@@
$S_{1,x}=\frac{m^{2}}{2}\!\!\!\sum_{a=1,2}\phi_{a,x}^{2}+\frac{\lambda}{4}\Big(\sum_{a=1,2}\phi_{a,x}^{2}\Big)^{2}\!\!+\!\!\sum_{a=1,2}J_{a}\phi_{a,x}$ (16)
@@
Due to the dependency of $S_{1,x}$ on the background shift $J_{a}$ in Eq. 16, each $\langle\phi_{1}\rangle_{J}$ calculation needs a training of a separate RBF network that depends on the actual value of the shift in order to be able to obtain the necessary $J(\phi_{1})$ function for the effective potential. The training of the RBF networks follows the same procedure that is described in the previous section. The determination of the $K$, $c_{k}$, $A$, and $a_{k}$ parameters is done in an automated way by training the networks for each $J$, using many different uniform grids for the centers, defined between $4,4]\times[-4,4]$, with different numbers of kernels in the range of $K\in[5^{2},30^{2}]$. For example, one configuration could consist of a uniform grid of centers in $1,1]\times[-1,1]$ with $K=5^{2}$, while another configuration could be a uniform grid in $2,2]\times[-2,2]$ with $K=10^{2}$. The width parameter is also changed in every network configuration in the range of $A\in[1,15]$. The $a_{k}$ parameters are determined with each of these parameters by solving the optimization problem in Eq. 6. Then, using randomly generated test data, the best model with the minimal mean squared error is chosen. Using this method, each $J$ could correspond to a different RBF network that, in a mean squared sense, optimally describes the given shifted functional forms.
@@
where $Z^{rbf}$ is the RBF approximated partition function. The $\mu$ independence is the consequence of severe cancellations due to the imaginary parts in the action, thus, in the phase quenched theory (i.e., when $\mathcal{B}_{p}=0$ in Eq. II) we do not expect this behavior. In contrast to the previous case, when we needed to introduce constant background shifts to the $S_{1,x}$ action in Eq. 16, now we only need to approximate one functional form given by the original form of $S_{1,x}$ shown in Eq. 4. The optimization procedure follows the previously described method, i.e., optimizing for $a_{k}$ using a set of $A$, $c_{k}$, and $K$ parameters, then choosing the model that best describes the test data in the mean squared sense. To show this in more detail, the training procedure has been carried out by assuming a two-dimensional uniform grid of center configuration, then fixing the number of kernels to $K=10^{2}$, and varying the widths parameters and the center configurations in the range of $A\in[0.5,3.1]$ and $c_{k}\in[-C,C]\times[-C,C]$, where the parameter $C$ determines the range of the centers and is chosen from $C\in[0.5,4.1]$. The determination of the center configurations this way means that we will have a $10\times 10$ uniform grid of centers with $\Delta c=2C/(K-1)$ resolution. To solve the optimization problem for the $a_{k}$ weight parameters, $2500$ input pairs $(\phi 1,\phi 2)$ were uniformly sampled from the domain $2,2]\times[-2,2]$, which was then used for training the corresponding networks. The determination of the best network configuration has been done by calculating the corresponding root mean squared values (RMSE) for the different RBF parametrizations using $1600$ test samples. The results can be followed in Fig. 6, where the RMSE dependence on $A$ and $C$ is shown on a contour plot, where the darkest parts correspond to the smaller errors. On the upper left subfigure the scatter plot of the true vs. predicted values is shown for the optimal case that corresponds to the parametrization $C=1.7$, $A=2.9$, while on the upper right scatter plot a ’bad’ parametrization is shown with $C=3.9$, $A=2.9$ that fails to reproduce the approximable function with a ’good’ accuracy. As the calculations for the silver blaze region only need one RBF network, we will use the previously determined optimal parameters with $K=10^{2}$ number of kernels, where the centers are distributed on a uniform grid in the range of $c_{k}\in[-1.7,1.7]\times[-1.7,1.7]$ with $\Delta c=3.4/9$ resolution.
Figure 6: RMSE dependence on the width and center parameterizations of the different RBF configurations with $K=10^{2}$ number of kernel functions. The minimum corresponds to the darkest value at $C=1.7$, and $A=2.9$. The top figures show the scatter plots of the true vs. estimated values for the optimum parametrization (left), and for another case, when the approximation is insufficient to describe the system (right).
By having the RBF approximated partition function, one could derive the closed-form expression for the expectation value of $\langle n\rangle$, or one could take the numerical derivatives by, e.g., a simple forward difference scheme. In this work we choose the latter with $\Delta\mu=0.001$. The calculations are done up until the critical chemical potential at a lattice size of $N^{4}=10^{4}$, and the results can be followed in Fig. 7, where the phase quenched values (i.e., when $\mathcal{B}_{p}=0$) are also shown for comparison.
@@
IV Numerical complexity
In this section the numerical and time complexity of the method and its dependence on dimensionality and lattice size will be discussed. In general, on a cubic lattice with $N$ number of lattice sites in each dimension, the complexity of the RBF method in momentum space can be described in $\mathcal{O}(KN^{D})$, where $K$ is the number of Gaussian kernels, $N$ is the number of lattice sites in one space-time direction, and $D$ is the dimensionality of the system. This general structure stays the same even if the number of lattice sites are different in the space and in the time directions, in which case one has $\mathcal{O}(KN_{space}^{D-1}N_{time})$. According to this, it is expected that the time it takes to calculate the different observables, such as the number density, scales linearly with the total number of lattice sites, which is very desirable when one wants to take, e.g., the thermodynamic or the continuum limits. Here, the time scaling will be discussed through the calculation of the number densities using the same system configuration that was used in the previous section, namely $m=1$ mass and $\lambda=1$ coupling strength.
To address the time complexity of the model, the number density defined in Eq. 19 is calculated at $N_{\mu}=24$ chemical potential values between $\mu\in[0,1.15]$ for the full theory, i.e., the same calculation is carried out as it is shown in Fig. 7. Algorithmically, the calculation can be separated into the following steps:
1) First, the parameters of the RBF network have to be optimized by the same method that is described in the previous section. This takes a fixed time $T_{RBF}$ that does not scale with the lattice size but only with the number of kernels.
2) The next step is the determination of $\ln Z^{rbf}$ for every chemical potential, which requires a D-dimensional nested loop (in this case $D=4$), where the $\mathcal{A}_{p}$ and $\mathcal{B}_{p}$ parameters defined in Eq. II have to be calculated at every momenta. As $\mathcal{A}_{p}$ also depends on the chemical potential, it cannot be taken out from the main loop, where we go through all $\mu$, thus, these nested loops have to be done $N_{\mu}$ times.
3) Lastly, the derivatives are calculated numerically by using a small $\Delta\mu$ step. This is done by making an extra calculation of $\ln Z^{rbf}$ for each $\mu$, so that the derivative can be approximated by using, e.g., the forward scheme as $(\ln Z^{rbf}(\mu+\Delta\mu)-\ln Z^{rbf}(\mu))/\Delta\mu$.
To summarize, the full calculation of the number density of the full theory using the RBF approximation on $N_{\mu}=24$ points needs the determination of the partition function $2N_{\mu}$ times, where in each calculation we have a 4-dimensional nested loop ($N^{4}$ terms) that represents the full sum over all momentum modes. The time it takes to do the full calculation will be addressed as follows:
$T=T_{RBF}+T_{L},$ (20)
where $T_{RBF}$ represents the time it takes to train one RBF network, while $T_{L}$ means the time to calculate the number densities at all the given chemical potentials. In general $T_{RBF}\ll T_{L}$, however its actual value depends on the optimization method, number of kernels, number of training/validation/test samples, etc. On a standard notebook (LG Gram, Intel Core i5, 16GB memory) the determination of the RBF parameters for one $(K,A,c_{k})$ configuration in the $K$ range we will do the comparisons ($K\sim 10^{2}$) is $T_{RBF}\sim 0.1$ [s]. Note that in the determination of the optimal configuration, e.g., in Fig. 6, one needs a set of different configurations, and the full optimization time will be larger.
Figure 8: Time complexity as a function of lattice size $N^{4}$ and number of Gaussian kernels $K$ in computing the number density at $N_{\mu}=24$ chemical potential values.
To examine the scaling behavior of $T_{L}$, the calculations have been made by setting $K=5^{2},10^{2},15^{2}$ and $N^{4}=2^{4},4^{4},6^{4},8^{4},10^{4},12^{4}$. The results can be seen in Fig. 8, where the assumed linear scaling with $K$ and $N^{4}$ is reproduced. According to the obtained results, the time that it takes to calculate the number densities in $24$ points (using the the optimal RBF network that is determined in Fig. 6) is $T_{L}\sim 14$ [s].
According to the algorithm, the numerically most expensive part of the calculations is the 4-dimensional nested loops, which has to be done in every step due to the $\mu$ dependence of the $\mathcal{A}_{p}$ and $\mathcal{B}_{p}$ parameters. The time it takes to calculate one point also depends on the number of kernel functions $K$, that is also an expected behavior of the model.
V Conclusions
@@
The RBF method is able to provide results in a very fast manner, e.g., the calculation regarding the silver blaze problem in Fig. 7 on a $10^{4}$ lattice took less than a minute on a standard notebook without any optimization or using any GPU or special hardware. Due to the generality of the model, there is a possibility to extend it to include fermionic fields and non-abelian gauge theories and address problems in QCD at finite densities as well. On the technical side, due to its versatility, the RBF model could also be extended or formulated in other ways that could be more suitable to specific problems.
VI acknowledgments
@@ Acknowledgements.
This work was supported by the Korea National Research Foundation under Grant No. 2023R1A2C300302311 and 2023K2A9A1A0609492411, and the Hungarian OTKA fund K138277.